OpenStudy (anonymous):

find the geometric mean of (1/4) and 80 if necessary give the answer in simplest radical form . please show work thank you

OpenStudy (anonymous):

well that was wrong \[80\times \frac{1}{4}=20\] \[\sqrt{20}=\sqrt{4\times 5}=2\sqrt{5}\]

OpenStudy (anonymous):

not that you should not ask, but is this clear? the steps i mean.

OpenStudy (anonymous):

yess they are and thank you but i can't find a scientific calculator online

Directrix (directrix):

(1/4)/x = x/ 80 x^2 = 20 x = 2 √ 5

Directrix (directrix):

http://www.calculator.com/calcs/calc_sci.html

OpenStudy (anonymous):

i bet you can divide 80 by 4 without a calculator (even though my first answer was wrong!)

OpenStudy (anonymous):

it was ( 1/4) not 4

OpenStudy (anonymous):

yes i know, but 80 times 1/4 is the same as 80 divided by 4

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

believe me i am not trying to be critical, it is just that if god forbid you have to take a test or get a quiz on this i hope it is clear what to do. multiply take the square root write in simplest radical form by looking for perfect squares in the radical

OpenStudy (anonymous):

most on line calculators will not write things in simplest radical form for you neither will hand held ones. you have to do it yourself, or use worlfamalpha

OpenStudy (anonymous):

that is if you use a calculator you will see \[\sqrt{20}=4.472...\] not \[\sqrt{20}=2\sqrt{5}\]

OpenStudy (anonymous):

yeaa thanks you